General Note on 1203SH CTD Casts & Processing
2017 update - CTD+Final bottle data remerged to generate 82-column CTD.csvs. All files replotted to QC. Datacode "9" added to Cast 029 ISUSQ, for missing data. Datacode "9" added to Cast 030 upcast from 138 - 0m.
1). Short delay in RV Shimada departure required dropping SCCOOS stations to
complete the 66 standard CalCOFI stations.
2). Final completed station talley was 57 out of 66 hydrographic stations
3). Shimada CTD does not have 3500+m of wire so no deep cast on 90.90 was
Cast 2: Bottle #2 fired but did not close - no water samples
Cast 3: Upcast profiles during the cast looked noisy.
Cast 4: Transmissometer noise from ~190m to 515m; cable will be serviced post-cast
Cast 8 - 11: Rough weather, big rolls
Cast 11: RAJ training on CTD ops begins with JLW; continues on casts 14 & 15; RAJ solos Cast 19
Cast 16: Extra bottle tripped at 62m (accidental); two extra LTER surface bottles tripped
Cast 18: Big rolls; deep chl max @ ~100m
Cast 23: Wire angle issues so CTD yo-yo'd a bit during the upcast to hit target depths
Cast 28: Very calm sea conditions
Cast 29: Santa Monica Basin; 440m bottle tripped as #4, CTD lowered back to 480m as #5, then to 320m as #6 (yo-yo'd); no ISUS data
Cast 30: No marker for bottle #20, since dup of #19 marker record cloned. Secondary conductivity/salinity flatlined on upcast from 138-0m.
Cast 32: Bottle #2 was closed but O2 temp is higher than expected, possible mistrip; carousel serviced post-cast
Cast 43: Santa Barbara Basin; shallow hypoxic zone; noisy fluorometer; moderate seas
Cast 48: Bottle #9 lid ajar, water samples compromised; bottle replaced
Cast 49: O2 on downcast 'meandering' - actual features, stratification
Cast 53: Rough conditions; ISUS was not powered, no ISUS data this cast; battery replaced post-cast
10Mar2017 update to 20-1203SH_CTDFinalQC.zip file:
Seasoft-generated .asc & .hdr files are available renamed as 20-1203SH_LLLLSSSS_###d or u.asc & .hdr. Voltages in the .asc files were also relabled. Since this makes it difficult to reprocess & merge with bottle data if necessary, the original .asc, .hdr, & .btl have also been archived under the "orginal_asc-hdr" subdirectory (in the archive). Their voltage header labels have not been changed so refer to the corresponding .hdr file for sensor configuration. This cruise did not have any sensors rearranged.
JRW 24Jun2014/ 10Mar2017